# Dial-In LLM: Human-Aligned LLM-in-the-loop Intent Clustering for Customer Service Dialogues

### v2

@@ Appendix B Prompt Template and More Results for LLM Utilities
Considering the resource constraints of industrial deployment, we evaluated the coherence performance using the OpenAI APIs: GPT-3.5, GPT-4, and GPT-4o, achieving accuracies of 89.58%, 93.54%, and 94.17%, respectively. These findings indicate that a smaller, fine-tuned LLM, such as qwen7b, can outperform these advanced models and potentially reduce the costs associated with API requests [^3] The fine-tuned LLMs for coherence evaluation and cluster naming will be released upon acceptance.. To highlight the importance of fine-tuning, we also tested a vanilla Qwen2.5 model without fine-tuning, which resulted in a significantly lower accuracy of 75.63%.
Model Sampling Method 1 2 3 4 5
qwen14b convex - 96.25% 95.42% 95.42% 95.21%
random (n=10) 95.83% 94.38% 93.33% 92.92% 92.92%
random (n=20) 94.58% 95.00% 94.38% 94.17% 94.17%
chatglm3-6b convex - 92.08% 90.83% 91.04% 90.42%
random (n=10) 92.29% 88.75% 85.83% 84.58% 85.21%
random (n=20) 90.42% 89.58% 89.38% 89.58% 88.98%

### v3

@@ Appendix B Prompt Template
Appendix C More Results: Evaluation with Proprietary LLM
Given the resource-intensive nature of LLM fine-tuning, we conducted further experiments in coherence evaluation using proprietary LLMs accessed through OpenAI APIs: GPT-3.5, GPT-4, and GPT-4o. These advanced and widely used models offer strong performance in various tasks but also come with higher operational costs due to token consumption. Using the same evaluation set of 480 clusters, these models achieved accuracies of 89.58%, 93.54%, and 94.17%, respectively. The results highlight that a smaller, fine-tuned LLM, such as Qwen-2.5-7b (with an accuracy of 96.25%), can outperform these advanced models while significantly reducing costs associated with API usage [^4] The fine-tuned Chinese LLMs for coherence evaluation and cluster naming will be released upon acceptance.. For comparison, a vanilla Qwen-2.5-7b model (without fine-tuning) was also evaluated, yielding a substantially lower accuracy of 75.63%, emphasizing the essence of fine-tuning.